# Singular dynamics for discrete weak K.A.M. solutions of exact twist maps

### v1

@@ 5. Singular dynamics on $\operatorname{Graph}(c+\nabla^{+}u_{c})$
Let $\Omega$ be a bounded convex open neighborhood of $x$. Since $S$ is locally semiconcave, we have $S(\cdot,y)$ is semiconcave on $\Omega$. By utilizing items (i) and (iv) of Proposition 2.2, there exists $C\geq 0$ such that
@@
Hence, $\partial_{1}S(x,y)\in\nabla^{+}(-u)(x)$ follows from item (ii) of Proposition 2.2. Since $\nabla^{+}(-u)(x)=-\nabla^{-}u(x)$, we conclude that $-\partial_{1}S(x,y)\in\nabla^{-}u(x)$.
Now, we move on to the second inclusion, $\partial_{2}S(x,y)\in\nabla^{+}T^{-}[u](y)$. According to the definition of the discrete backward Lax-Oleinik operator $T^{-}$, we have $T^{-}[u](\cdot)\leq u(x)+S(x,\cdot)$. Hence it follows from items (iii) and (iv) of Proposition 2.2 that
@@
The ferromagnetic condition of $S$ asserts that if $x_{1}\neq x_{2}$, then $\partial_{2}S(x_{1},y)\neq\partial_{2}S(x_{2},y)$. Hence $\nabla^{+}T^{-}[u](y)$ is not a singleton set. It follows from item (iv) of Proposition 2.2 that $T^{-}[u]$ is not differentiable at $y$.
Thanks to item (c) of Remark 2.4, we only need to demonstrate that if $x$ is the unique point such that $T^{-}[u](y)=u(x)+S(x,y)$, then $T^{-}[u]$ is differentiable at $y$.
@@
Hence $\partial_{2}S(x,y)\in\nabla^{-}T^{-}[u](y)$ by the definition of superdifferential. Using item (a) of Lemma 2.5 and item (iv) of Proposition 2.2, we have that $T^{-}[u]$ is differentiable at $y$.
@@
Proof of Lemma 2.12.
(a). If $(x,y)\in\mathcal{NS}(T^{-}[u])$, then by the definition, we have
$T^{-}[u](y)-T^{-}[u](x)=S(x,y)-\bar{S}.$
That is,
$\left[T^{-}[u](y)-u(x)-S(x,y)\right]+\left[u(x)-T^{-}[u](x)+\bar{S}\right]=0.$
From the definitions of $T^{-}$ and sub-actions, we see that
$T^{-}[u](y)-u(x)-S(x,y)\leq 0,\ \quad\ u(x)-T^{-}[u](x)+\bar{S}\leq 0$
respectively. Therefore,
$T^{-}[u](y)-u(x)-S(x,y)=0,\ \quad\ u(x)-T^{-}[u](x)+\bar{S}=0.$
Using (a) of Lemma 2.5, since $T^{-}[u](y)-u(x)-S(x,y)=0$, we have that
$-\partial_{1}S(x,y)\in\nabla^{-}u(x).$
By (c) of Remark 2.4, one can find $w\in\mathbb{R}^{d}$ such that $T^{-}[u](x)=u(w)+S(w,x).$ Using (a) of Lemma 2.5, we know that
$\partial_{2}S(w,x)\in\nabla^{+}T^{-}[u](x).$
Since $u$ is a sub-action, i.e., $u+\bar{S}\leq T^{-}[u]$, and $u(x)+\bar{S}=T^{-}[u](x)$, it follows from the item (iii) of Proposition 2.2 that
$\nabla^{+}u(x)\supset\nabla^{+}T^{-}[u](x)\quad\text{and}\quad\nabla^{-}u(x)\subset\nabla^{-}T^{-}[u](x).$
In conclusion, we have
$-\partial_{1}S(x,y)\in\nabla^{-}u(x)\subset\nabla^{-}T^{-}[u](x);$
$\partial_{2}S(w,x)\in\nabla^{+}T^{-}[u](x)\subset\nabla^{+}u(x).$
By (iv) of Proposition 2.2, we claim that both $u$ and $T^{-}[u]$ are differentiable at $x$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)=dT^{-}[u](x)=-\partial_{1}S(x,y)$.

### v2

@@ Appendix B Twist maps of the $2$-dimensional annulus
Let $\Omega$ be a bounded convex open neighborhood of $x$. Since $S$ is locally semiconcave, we have $S(\cdot,y)$ is semiconcave on $\Omega$. By utilizing item (i) of Proposition A.2, there exists $C\geq 0$ such that
@@
Hence, $\partial_{1}S(x,y)\in\nabla^{+}(-u)(x)$ follows from item (i) of Proposition A.2. Since $\nabla^{+}(-u)(x)=-\nabla^{-}u(x)$, we conclude that $-\partial_{1}S(x,y)\in\nabla^{-}u(x)$.
Now, we move on to the second inclusion, $\partial_{2}S(x,y)\in\nabla^{+}T^{-}[u](y)$. According to the definition of the discrete backward Lax-Oleinik operator $T^{-}$, we have $T^{-}[u](\cdot)\leq u(x)+S(x,\cdot)$. Hence it follows from items (ii) and (iii) of Proposition A.2 that
@@
The ferromagnetic condition of $S$ asserts that if $x_{1}\neq x_{2}$, then $\partial_{2}S(x_{1},y)\neq\partial_{2}S(x_{2},y)$. Hence $\nabla^{+}T^{-}[u](y)$ is not a singleton set. It follows from item (iii) of Proposition A.2 that $T^{-}[u]$ is not differentiable at $y$.
Thanks to item (a) of Remark 2.4, we only need to demonstrate that if $x$ is the unique point such that $T^{-}[u](y)=u(x)+S(x,y)$, then $T^{-}[u]$ is differentiable at $y$.
@@
Hence $\partial_{2}S(x,y)\in\nabla^{-}T^{-}[u](y)$ by the definition of superdifferential. Using item (a) of Lemma 2.19 and item (iii) of Proposition A.2, we have that $T^{-}[u]$ is differentiable at $y$.
@@
Acknowledgments
The authors thank Prof. P. Thieullen for a careful reading of the manuscripts and many helpful suggestions during his visit at Beijing Normal University.